OpenStudy (anonymous):

Solve. x2 + 8x – 8 = 0 A. 05_14_UT_4a.gif B. 05_14_UT_4c.gif C. 05_14_UT_4b.gif D. 05_14_UT_4d.gif

OpenStudy (anonymous):

sorry

OpenStudy (solomonzelman):

Whatever your answer choices are... do you prefer quadratic formula or completing the square ?

OpenStudy (anonymous):

quadratic formula

OpenStudy (solomonzelman):

Yes, so \(\LARGE\color{blue}{\bf x=~ \frac{-b±\sqrt{b^{2}-4(a)(c)}}{2(a)} }\) can you plug in numbers ?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \frac{-(8)±\sqrt{(8)^{2}-4(1)(-8)} }{2(1)} }\) \(\LARGE\color{blue}{ \bf \frac{-(8)±\sqrt{64-(-32)} }{2} }\) \(\LARGE\color{blue}{ \bf \frac{-8±\sqrt{96} }{2} }\) \(\LARGE\color{blue}{ \bf \frac{-8±\sqrt{16 \times 6} }{2} }\) \(\LARGE\color{blue}{ \bf \frac{-8±4\sqrt{ 6} }{2} }\) \(\LARGE\color{blue}{ \bf -4±2\sqrt{ 6} }\)